OpenStudy (anonymous):

During the first part of a trip, a canoeist travels 80 miles at a certain speed. The canoeist travels 17 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 2 hours. What is the speed on the first and second part of the trip? (type an integer or a decimal. Round to the nearest hundredth)

OpenStudy (anonymous):

distance/speed = time First part of the trip: Distance is x and speed is 80: x/80 Second part of the trip: Distance is 17 and speed is (x-5): 17/(x-5) (x/80)+(17/(x-5))=2

OpenStudy (anonymous):

the mph for the first and second trip is how much b/c I see 17 but got lost

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

your saying the MPH for part 1 is 80 & part 2 is 17?

OpenStudy (anonymous):

that is distance not speed

OpenStudy (anonymous):

his equation is correct except its 80/x for the first part

OpenStudy (anonymous):

right I am trying to figure the speed in mph

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

Nope, distance/speed = time. Distance is on top and speed is on bottom. 17 is the distance on the second part. Ah yeah, had a typo on my equation, sorry.

OpenStudy (anonymous):

(80/x)+(17/(x-5))=2

OpenStudy (anonymous):

grr

OpenStudy (anonymous):

(x/80)+(17/(x-5))=2 is correct

OpenStudy (anonymous):

gah, no it's not, hahaha

OpenStudy (anonymous):

(80/x)+(17/(x-5))=2

OpenStudy (anonymous):

so the first part is 2?

OpenStudy (anonymous):

first part is x and second speed is x-5

OpenStudy (anonymous):

you have to solve equation to get values

OpenStudy (anonymous):

the first part for the speed mph being x makes no sense to me sorry

OpenStudy (anonymous):

Solve the whole equation altogether.

OpenStudy (anonymous):

No problem, what about that makes it sound weird to you?

OpenStudy (anonymous):

because it is a letter & not a number

OpenStudy (anonymous):

if t represents the time it takes to complete the first part of the trip then the speed of the of the first part of the trip is 80/t. The speed of the second part of the trip is 17/2+t, which equals 80/t -5. Solve for t and then find the speed.

OpenStudy (anonymous):

The correct answer for the first part is 49.46

OpenStudy (anonymous):

My equation gives you 49.46.

OpenStudy (anonymous):

the second is 44.46

OpenStudy (anonymous):

Correct, because we solve for 49.46. Then we also know the second part of the trip was 5 mph slower so 49.46-5=44.46.

OpenStudy (anonymous):

well that isn't the answer you provided lol

OpenStudy (anonymous):

I provided you with an equation to solve. If you solved my equation you would have ended up with the correct answer.

OpenStudy (anonymous):

The problem is that with your method we have two variables one is t and one is speed. It's a lot harder to solve problems with two variables and often you can't even do so. There is no need to use t as a variable because they give you t. It's 2 hours.